# Can AI Dream of Unseen Galaxies? Conditional Diffusion Model for Galaxy Morphology Augmentation

### v1

@@ Abstract
Observational astronomy relies on visual feature identification to detect critical astrophysical phenomena. While machine learning (ML) increasingly automates this process, models often struggle with generalization in large-scale surveys due to the limited representativeness of labeled datasets—whether from simulations or human annotation—a challenge pronounced for rare yet scientifically valuable objects. To address this, we propose a conditional diffusion model to synthesize realistic galaxy images for augmenting ML training data. Leveraging the Galaxy Zoo 2 dataset which contains visual feature – galaxy image pairs from volunteer annotation, we demonstrate that our model generates diverse, high-fidelity galaxy images closely adhere to the specified morphological feature conditions. Moreover, this model enables generative extrapolation to project well-annotated data into unseen domains and advancing rare object detection. Integrating synthesized images into ML pipelines improves performance in standard morphology classification, boosting completeness and purity by up to 30% across key metrics. For rare object detection, using early-type galaxies with prominent dust lane features ( $\sim$0.1in GZ2 dataset) as a test case, our approach doubled the number of detected instances—from 352 to 872—compared to previous studies based on visual inspection. This study highlights the power of generative models to bridge gaps between scarce labeled data and the vast, uncharted parameter space of observational astronomy and sheds insight for future astrophysical foundation model developments. Our project homepage is available at https://galaxysd-webpage.streamlit.app/.
@@
Submitted to AAS journals
1 Introduction
Modern astronomy increasingly relies on accurate and robust machine learning (ML) systems to efficiently process vast observational datasets (e.g., Ball et al., 2006; Speagle et al., 2019b; Ting et al., 2019; Wang et al., 2022; Robertson et al., 2023; Li et al., 2024). In astrophysical image analysis, ML techniques have found widespread applications across diverse scientific tasks, including galaxy morphology classification (e.g., Banerji et al., 2010; Kim & Brunner, 2017; Walmsley et al., 2023a; Vavilova et al., 2021), detection of rare astronomical phenomena such as galaxy mergers, strong gravitational lensing systems, and ultra-diffuse galaxies (e.g., Li et al., 2020; Omori et al., 2023; Stein et al., 2022; Keerthi-Vasan et al., 2023; Thuruthipilly et al., 2025), artifact detection and removal (e.g., Zhang & Bloom, 2020; Xu et al., 2023; Tanoglidis et al., 2022; Zhang & Brandt, 2021), and image restoration through super-resolution and denoising (e.g., Sweere et al., 2022; Dabbech et al., 2022; Terris et al., 2023; Vavilova et al., 2021; Liu et al., 2025).
@@ 1 Introduction
To increase the generalizability of ML models in research environments, various attempts have been made. For example, unsupervised/self-supervised learning (e.g., Bishop, 2006; Goodfellow et al., 2016; Gui et al., 2023) leverages unlabeled data to learn useful representations without the need for extensive labeled datasets. This approach has been applied in astronomy to improve the performance of models on tasks such as galaxy morphology classification and anomaly detection (e.g., Hayat et al., 2021; Wei et al., 2022; Walmsley et al., 2022; Desmons et al., 2024; Mohale & Lochner, 2024; Riggi et al., 2024) and further inspired the recent development of astronomical foundational models (e.g., Walmsley et al., 2022; Sun et al., 2023b; Leung & Bovy, 2024; Walmsley et al., 2024).
Semi-supervised learning (e.g., van Engelen & Hoos, 2020; Yang et al., 2021) combines a small amount of labeled data with a large amount of unlabeled data during training, which has been shown to enhance model performance in scenarios where labeled data is scarce (e.g., Ćiprijanović et al., 2023; Slijepcevic et al., 2022). Transfer learning (e.g., Zhuang et al., 2019) uses pre-trained models from related domains and performs domain adaptation for specific tasks for deployments, boosting astrophysical ML model deployments in real-world research (e.g., Kim et al., 2021; O’Briain et al., 2021; Cavuoti et al., 2024).
Active learning (e.g., Zhan et al., 2022) iteratively selects the most informative samples for labeling, thereby reducing the labeling effort while maximizing model performance, which has been employed to label rare objects in astronomy (e.g., Leoni et al., 2022; Pruzhinskaya et al., 2023; Andersson et al., 2024). Reinforcement learning Sutton & Barto (2018) learns from self-play experiences guided by scientific rewards, and while underutilized in astronomy, shows promise for agent-based discovery in astrophysical ML development (e.g., Tenachi et al., 2023; Sun et al., 2024b).
While the methodologies mentioned above approach the challenge of data scarcity from different angles, recent advancements in deep generative models, particularly diffusion-based models for image and video generation (e.g., Dhariwal & Nichol, 2021; Rombach et al., 2021; Ho et al., 2022; Gupta et al., 2023; Ma et al., 2025), open new possibilities for addressing this issue. Today’s generative models not only produce high-quality, realistic images but also enable generative extrapolation into unseen domain. For instance, in the automotive industry, diffusion models can simulate various driving scenarios that may rarely occur in real-world data, such as extreme weather conditions, vehicle malfunctions, or unexpected pedestrian behaviors (e.g., Gao et al., 2024; NVIDIA et al., 2025).
By extrapolating on those rare scenarios, manufacturers can train autonomous driving systems more comprehensively, improving their ability to handle rare and critical situations, thus enhancing safety and reducing the costs associated with extensive real-world testing. As a result, research outside astronomy has begun to explore these models as potential world simulator to reduce the costs associated with robotic training, autonomous driving, and computer vision tasks (e.g., Zhu et al., 2023; Azizi et al., 2023; Song et al., 2023).
@@
In this study, we demonstrate the capabilities of our generative model, which is trained on Galaxy Zoo 2, one of the largest citizen science galaxy morphological dataset to date. Our generative model could learn and generate diverse, high-fidelity galaxy images that accurately capture the visual characteristics of real astronomical objects and adhere closely to specified morphological conditions. By integrating these synthetic images into machine learning workflows, we observe substantial enhancements in model performance within real-world astronomical research contexts, especially for the analysis of rare and scientifically interesting celestial objects.
@@
We validate our approach through two key tasks: classical galaxy morphology classification and rare object detection. In the realm of classical morphology classification, which encompasses a series of binary classification tasks such as differentiating early- from late-type galaxies and determining the presence of bulge/disk/bar, our model achieves up to a 30% improvement in both completeness and purity metrics. For the rare object detection task, we focus on identifying early-type galaxies with prominent dust lane features. Given that previous studies Kaviraj et al. (2012); Shabala et al. (2012); Davis et al. (2015) have simply identified 352 positive samples (approximately 0.1% of the GZ2 dataset), our approach leveraging those synthetic early-type dust-lane galaxies successfully uncovers an additional 520 positive samples that were overlooked due to the limitations of visual inspection in archival data.
@@
This work demonstrates the potential of generative models as surrogates to address data scarcity in astronomy, enhancing machine learning applications for large - scale surveys and promoting the development of astrophysical image foundation models. The paper is structured as follows: Section 2 details the diffusion model, training, and inference; Section 3 describes the Galaxy Zoo 2 dataset and data preprocessing; Section 4 shows the model’s ability to generate high - fidelity galaxy images matching specified morphological features and presents its applications in classical galaxy morphology classification and the detection of rare early - type galaxies with dust lanes (about 0.1% of the dataset). Section 5 explores broader scientific applications beyond galaxy morphology and synergies with other algorithms before concluding.
All models, data, and code used in this study are publicly accessible in https://github.com/chenruiRae/GalaxySD. All machine learning models in this study are trained in one NVIDIA GeForce RTX 3090 Graph Processing Unit. Where necessary, we assume parameters for the flat $\Lambda$CDM cosmology determined by the Planck Collaboration et al. (2020) (i.e., $\mathrm{h}=0.674$, $\Omega_{\mathrm{m}}=0.315$).
@@
Figure 1: Schematic diagram of our diffusion model and its applications to downstream tasks. For the training processes shown in the left two boxes, the training set consists of real images and descriptive morphological prompts (see Table 1 for more examples) from Galaxy Zoo 2, which are used to fine-tuning diffusion model. For inference, given morphological prompts, our fine-tuned model generates realistic and diverse galaxy images as demanded. These vast simulated galaxy images integrate with real data could be a high-quality augmented dataset used for other downstream ML tasks as detailed in Section 4.
2 Method
@@
Mathematically, given a input galaxy image $\mathbf{x_{0}}$ with corresponding morphological condition $\mathbf{c}$ (see Table 1 for examples), a conditional diffusion model tries to extend model the probability distribution $p(\mathbf{x_{0}}|\mathbf{c})$ by iteratively denoising a normally distributed variable $\mathbf{x}_{T}$, which can be viewed as modeling a reverse process of a Markov Chain $\mathbf{x}_{T},\mathbf{x}_{T-1},\dots,\mathbf{x}_{0}$. The reverse process is parameterized by a neural network with parameters $\theta$ and aims to predict a denoised variant $\mathbf{\epsilon}_{\theta}(\mathbf{x}_{t},\mathbf{c})$ from their input $\mathbf{x}_{t}$, where $\mathbf{x}_{t}$ is a noisy version of input $\mathbf{x}_{0}$. The training objective can be simplified as regressing the noise at each time step: $\mathrm{loss}=\mathbb{E}_{\mathbf{x},\mathbf{\epsilon}\sim\mathcal{N}(0,\mathbf{I}),t}\|\mathbf{\epsilon}-\epsilon_{\theta}(\mathbf{x}_{t},\mathbf{c})\|$.
@@
In practical implementation, we utilize the pre-trained diffusion model Stable-Diffusion-v1.5111https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5Rombach et al. (2021). This model adopts a U-Net architecture Ronneberger et al. (2015) combined with a text encoder. Input morphological conditions are first vectorized by the text encoder, then integrated into the image generation process via a cross-attention mechanism, detailed in Appendix C. These design features enable efficient and flexible processing of conditional inputs. We fine-tune the entire model using morphology condition-galaxy image pairs extracted from the GZ2 datasets (see Section 3 for details). Through full-parameter fine-tuning, the model adapts to the unique characteristics of astronomical imagery, enabling it to generate realistic, high-fidelity galaxy images that accurately capture subtle morphological details.
@@
We have observed that images generated solely with text conditioning consistently display a pattern of systematic over-brightness. This phenomenon stems from the singularities present in the diffusion process at the $t=1$ timestep, as documented in previous works (e.g., Lin et al., 2023a; Zhang et al., 2024). To address this issue, we incorporated randomly selected reference images and initiated the denoising process from the noised reference image. This strategic modification serves to effectively regulate the brightness levels of the generated images while preserving their visual fidelity.
@@
When performing generative extrapolation on unseen classes—specifically early-type galaxies with prominent dust lanes in this study, we observe that underrepresented features (e.g., dust lanes) can be diluted to varying degrees, depending on their prevalence in the training data. To address this, we implement a targeted feature enhancement strategy during inference. After converting input morphological prompts or descriptions into contextual embeddings, we scale the weights of the target features (e.g., dust lanes) to bias the cross-attention mechanism toward prioritizing these attributes. This weighting scheme enables the model to focus more intentionally on underrepresented characteristics, thereby producing synthetic images with enhanced fidelity to the specified features—an approach particularly critical when extrapolating beyond the distribution of the training dataset.
@@
After fully training the conditional generative model, our diffusion model learns the mapping between morphological conditions and corresponding visual features. It can synthesize galaxies from our dataset and extrapolate to rare or even unseen scenarios, for example, imagining an early-type galaxy exhibiting star-forming characteristics. These hypothetical entities, though rare or even never observed, hold scientific value. Incorporating these synthesized images into downstream machine learning model training pipelines can notably improve model performance on ill-defined problems and may lead to potential unique scientific discoveries.
3 Data

### v2

@@ Abstract
Observational astronomy relies on visual feature identification to detect critical astrophysical phenomena. While machine learning (ML) increasingly automates this process, models often struggle with generalization in large-scale surveys due to the limited representativeness of labeled datasets—whether from simulations or human annotation—a challenge pronounced for rare yet scientifically valuable objects. To address this, we propose a conditional diffusion model to synthesize realistic galaxy images for augmenting ML training data (hereafter GalaxySD). Leveraging the Galaxy Zoo 2 dataset which contains visual feature – galaxy image pairs from volunteer annotation, we demonstrate that GalaxySD generates diverse, high-fidelity galaxy images that closely adhere to the specified morphological feature conditions. Moreover, this model enables generative extrapolation to project well-annotated data into unseen domains and advancing rare object detection. Integrating synthesized images into ML pipelines improves performance in standard morphology classification, boosting completeness and purity by up to 30% across key metrics. For rare object detection, using early-type galaxies with prominent dust lane features ( $\sim$0.1% in GZ2 dataset) as a test case, our approach doubled the number of detected instances—from 352 to 872—compared to previous studies based on visual inspection. This study highlights the power of generative models to bridge gaps between scarce labeled data and the vast, uncharted parameter space of observational astronomy and sheds insight for future astrophysical foundation model developments. Our project homepage is available at https://galaxysd-webpage.streamlit.app/.
@@
I Introduction
Modern astronomy increasingly relies on accurate and robust machine learning (ML) systems to efficiently process vast observational datasets (e.g., Ball et al., 2006; Speagle et al., 2019b; Ting et al., 2019; Wang et al., 2022; Robertson et al., 2023; Li et al., 2024). In astrophysical image analysis, ML techniques have found widespread applications across diverse scientific tasks, including galaxy morphology classification (e.g., Banerji et al., 2010; Kim & Brunner, 2017; Walmsley et al., 2023a; Vavilova et al., 2021), detection of rare astronomical phenomena such as galaxy mergers, strong gravitational lensing systems, and ultra-diffuse galaxies (e.g., Li et al., 2020; Omori et al., 2023; Stein et al., 2022; Keerthi-Vasan et al., 2023; Thuruthipilly et al., 2025), artifact detection and removal (e.g., Zhang & Bloom, 2020; Xu et al., 2023; Tanoglidis et al., 2022; Zhang & Brandt, 2021), and image restoration through super-resolution and denoising (e.g., Sweere et al., 2022; Dabbech et al., 2022; Terris et al., 2023; Vavilova et al., 2021; Liu et al., 2025a).
@@ I Introduction
To increase the generalizability of ML models in research environments, various attempts have been explored beyond traditional supervised learning. Methods such as unsupervised, semi-supervised, transfer, active and reinforcement learning have been explored to mitigate data scarcity in astronomy (e.g., Hayat et al., 2021; Walmsley et al., 2022; Slijepcevic et al., 2022; O’Briain et al., 2021; Leoni et al., 2022; Tenachi et al., 2023; Sun et al., 2024b). These approaches leverage unlabeled data, pretrained models, selective labeling or reward mechanism to enhance performance in low-data regimes.
Although the methodologies mentioned above approach the challenge of data scarcity from different angles, recent advances in deep generative models, particularly diffusion-based models for image and video generation (e.g., Dhariwal & Nichol, 2021; Rombach et al., 2021; Ho et al., 2022; Gupta et al., 2023; Ma et al., 2025), offer a powerful new approach to alleviate data scarcity. Beyond generating high-quality realistic samples, diffusion models enable extrapolation into rare or unseen scenarios, making them valuable as world simulators. In computer vision and automation, they help simulate various conditions, including rare and critical situations, reducing the dependency on costly real-world data collection and improving the robustness of the model (e.g., Zhu et al., 2023; Azizi et al., 2023; Song et al., 2023).
@@
In this study, we demonstrate the capabilities of our generative model, GalaxySD, which is trained on Galaxy Zoo 2, one of the largest citizen science galaxy morphological datasets to date. The generative model could learn and generate diverse, high-fidelity galaxy images that accurately capture the visual characteristics of real astronomical objects and adhere closely to specified morphological conditions. By integrating these synthetic images into machine learning workflows, we observe substantial enhancements in model performance within real-world astronomical research contexts, especially for the analysis of rare and scientifically interesting celestial objects.
@@
We validate our approach through two key tasks: classical galaxy morphology classification and rare object detection. In the realm of classical morphology classification, which encompasses a series of binary classification tasks such as differentiating early- from late-type galaxies and determining the presence of spiral/disk/bar/bulge, classifier trained on dataset augmented by GalaxySD (hereafter, synthetic-augmented classifier) achieves up to a 30% improvement in both completeness and purity metrics. For the rare object detection task, we focus on identifying early-type galaxies with prominent dust lane features. Given that previous studies Kaviraj et al. (2012); Shabala et al. (2012); Davis et al. (2015) have simply identified 352 positive samples (approximately 0.1% of the GZ2 dataset), our approach leveraging those synthetic early-type dust-lane galaxies successfully uncovers an additional 520 positive samples that were overlooked due to the limitations of visual inspection in archival data.
@@
This work demonstrates the potential of generative models as surrogates to address data scarcity in astronomy, enhancing machine learning applications for large - scale surveys and promoting the development of astrophysical image foundation models. The paper is structured as follows: Section II details the diffusion model, training, and inference; Section III describes the Galaxy Zoo 2 dataset and data preprocessing; Section IV shows the model’s ability to generate high - fidelity galaxy images matching specified morphological features and presents its applications in classical galaxy morphology classification and the detection of rare early - type galaxies with dust lanes (about 0.1% of the dataset). Section V explores broader scientific applications beyond galaxy morphology and synergies with other algorithms before concluding.
[…]
@@
Figure 1: Schematic diagram of GalaxySD, evaluation of synthetic images, and applications to downstream tasks. For the training processes shown in the upper panel, the training set consists of real galaxy images and descriptive morphological labels (refer to Figure 2) from Galaxy Zoo 2, which are used to fine-tuning (abbreviated as FT) diffusion model. In the diffusion model, $x_{0}$} denotes the original galaxy image, $x_{T}$} the noised image sampled from a normal distribution, and $x_{0}$$^{\prime}$ the reconstructed galaxy image predicted by the neural network conditioned on $x_{T}$} and the input text. For inference, given morphological prompts (refer to Table 1), our fine-tuned model, GalaxySD generates realistic and diverse galaxy images as demanded. These generated images are evaluated by three metrics-realism, diversity and consistency as described in Section IV.1. In practical applications, these simulated galaxy images, integrated with limited real data, could form a high-quality augmented dataset for training a binary classifier, achieving better performance (we use an upward arrow to denote this) than other ML methods and found new galaxies in Galaxy Zoo SDSS, as detailed in Sections IV.2 and IV.3.
II Method
II.1 Diffusion Model
@@
Mathematically, given an input galaxy image $\mathbf{x_{0}}$ with corresponding morphological condition $\mathbf{c}$ (see Table 1 for examples), a conditional diffusion model tries to extend the model the probability distribution $p(\mathbf{x_{0}}|\mathbf{c})$ by iteratively denoising a normally distributed variable $\mathbf{x}_{T}$, which can be viewed as modeling a reverse process of a Markov chain $\mathbf{x}_{T},\mathbf{x}_{T-1},\dots,\mathbf{x}_{t},\dots,\mathbf{x}_{0}$. The reverse process is parameterized by a neural network with parameters $\theta$ and aims to predict a denoised variant $\mathbf{\epsilon}_{\theta}(\mathbf{x}_{T},\mathbf{c})$ from their input $\mathbf{x}_{T}$, where $\mathbf{x}_{T}$ is a noisy version of input $\mathbf{x}_{0}$. The training objective can be simplified as regressing the noise at each time step $t$: $\mathrm{loss}=\mathbb{E}_{\mathbf{x},\mathbf{\epsilon}\sim\mathcal{N}(0,\mathbf{I}),t}\|\mathbf{\epsilon}-\epsilon_{\theta}(\mathbf{x}_{t},\mathbf{c})\|$. Finally, after the reverse denoising process, the reconstructed galaxy image $\mathbf{x}_{0}^{\prime}$ serves as an approximation of the original input $\mathbf{x}_{0}$.
II.2 Cross-attention Mechanism in Conditional Diffusion Model
In the diffusion model, conditioning on textual prompts is achieved through a cross-attention mechanism integrated into the U-Net backbone of the denoising process. Instead of conditioning the model via concatenation or direct feature injection, it employs text-image cross-attention mechanism at multiple layers, enabling semantically aligned image generation guided by prompts.
Let $\mathbf{x}_{t}\in\mathbb{R}^{C\times H\times W}$ denote the latent visual features at denoising timestep $t$, where $C$ is the channel, the characteristic dimension of the image encoder and $H\times W$ is the spatial resolution. Then let $\boldsymbol{\tau}=[\tau_{1},\ldots,\tau_{n}]$, with each $\tau_{i}\in\mathbb{R}^{d}$, denote the embedded text tokens from a frozen text encoder, where $d$ is the text embedding dimension. To enable cross-attention, the image features are projected from $\mathbb{R}^{C}$ to $\mathbb{R}^{d}$ at each spatial location via a learned linear mapping $\psi:\mathbb{R}^{C}\rightarrow\mathbb{R}^{d}$, aligning them with the text feature space. This yields a query matrix $\mathbf{Q}\in\mathbb{R}^{(HW)\times d}$. The keys and values, $\mathbf{K},\mathbf{V}\in\mathbb{R}^{n\times d}$, are derived from the text embeddings. The attention output is computed as,
$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)\mathbf{V}$ (1)
The attention output is a matrix of size $\mathbb{R}^{(HW)\times d}$, producing a text-conditioned representation for each spatial location in the visual feature map. The attention value is reshaped to $\mathbb{R}^{H\times W\times d}$ to match the spatial layout. Since the original feature map has channel dimension $C$, the output is further projected back to $\mathbb{R}^{C}$ by an inverse linear layer $\rho:\mathbb{R}^{d}\rightarrow\mathbb{R}^{C}$ before being added to the input.
Thus, the final output of the ResNet block with cross-attention is:
$\mathbf{x}_{\mathrm{out}}=\mathbf{x}_{\mathrm{in}}+\rho\left(\mathrm{Attention}(\psi(\mathbf{x}_{\mathrm{in}}),\boldsymbol{\tau},\boldsymbol{\tau})\right)$ (2)
Here, $\psi$ and $\rho$ are learned linear projections that ensure the query vector matches the key/value dimensions, so that the attention output can be fused back with the input via a residual connection. This makes the whole operation dimensionally consistent with $\mathbf{x}_{\mathrm{in}}\in\mathbb{R}^{C\times H\times W}$.
Unlike standard self-attention mechanism where $\mathbf{Q},\mathbf{K},\mathbf{V}$ are all derived from the same input, cross-attention distinguishes between the source of queries from images and the keys, values from text. This enables spatially aware modulation of visual features based on textual semantics. Also, multi-head attention Vaswani et al. (2017) is employed within the cross-attention blocks to capture long-range dependencies and diverse semantic relationships between image regions and textual concepts. It operates by computing multiple attention functions in parallel, each attending to different subspaces of the input, and then concatenating their outputs to yield a refined alignment between visual and textual features.
During inference, to enhance specific morphological features such as dust lanes, we apply weighted prompts by scaling the corresponding token embeddings, using a syntax like {dust lane:1.3}. Formally, let $\mathbf{E}=[\mathbf{e}_{1},\dots,\mathbf{e}_{n}]^{\top}\in\mathbb{R}^{n\times d^{\prime}}$ be the matrix of token embeddings, where $\mathbf{e}_{j}\in\mathbb{R}^{d^{\prime}}$ is the embedding for the $j$-th token. The key and value matrices in the cross-attention module are obtained by linear projection:
$\mathbf{K}=\mathbf{E}\cdot\mathbf{W}_{K},\quad\mathbf{V}=\mathbf{E}\cdot\mathbf{W}_{V}$ (3)
where $\mathbf{W}_{K},\mathbf{W}_{V}\in\mathbb{R}^{d\times d^{\prime}}$ are learnable projection matrices. To emphasize certain concepts, we modify the embeddings by applying a scalar weight $w_{j}>1$ to the corresponding token embedding:
$\mathbf{e}_{j}^{\prime}=w_{j}\cdot\mathbf{e}_{j}$ (4)
This leads to increased magnitudes in both $\mathbf{K}$ and $\mathbf{V}$, which amplifies attention logs and strengthens the contribution of that token to the attended representation. The resulting softmax distribution becomes biased toward the emphasized feature, thus guiding the model to generate images with enhanced visual presence of the specified concept. In practice, we find weights in the range of 1.1–1.4 to be effective without overpowering the rest of the prompt semantics. A detailed ablation study validating this weight selection is provided in Appendix C.
II.3 Implementation of GalaxySD
@@
In practical implementation, we utilize the pretrained diffusion model Stable-Diffusion-v1.5111https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5Rombach et al. (2021). This model adopts a U-Net architecture Ronneberger et al. (2015) combined with a text encoder. The input morphological conditions are first vectorized by the text encoder, and then integrated into the image generation process via a cross-attention mechanism. These design features enable efficient and flexible processing of conditional inputs. We performed full-parameter fine-tuning, including both the U-Net backbone parameters, which are optimized to better capture galaxy morphology, and the text-to-vector embedding layers of the text encoder, which are updated so that morphological condition descriptions are more effectively mapped to visual features. In this way, both the image generation backbone and the conditioning layers are jointly adapted using the morphology condition–galaxy image pairs from the GZ2 dataset (see Section III), allowing the model to generate high-fidelity galaxy images that accurately reflect subtle morphological details.
@@
We have observed that images generated solely with text conditioning consistently display a pattern of systematic over-brightness in the resulting RGB images. This refers to a visually perceptual overexposure effect that galaxy images appear significantly brighter or washed out in color, rather than a photometric deviation in integrated flux. This phenomenon stems from the singularities present in the diffusion process at the $t=1$ timestep, as documented in previous works (e.g., Lin et al., 2023a; Zhang et al., 2024). To address this issue, we incorporated randomly selected reference images and initiated the denoising process from the noised reference image. This strategic modification serves to effectively regulate the brightness levels of the generated images while preserving their visual fidelity.
@@
When performing generative extrapolation on unseen classes, specifically early-type galaxies with prominent dust lanes in this study, we observe that underrepresented features (e.g., dust lanes) can be diluted to varying degrees, depending on their prevalence in the training data. To address this, we implement a targeted feature enhancement strategy during inference. After converting input morphological prompts or descriptions into contextual embeddings, we scale the weights of the target features (e.g., dust lanes) to bias the cross-attention mechanism toward prioritizing these attributes. This weighting scheme enables the model to focus more intentionally on underrepresented characteristics, thereby producing synthetic images with enhanced fidelity to the specified features—an approach particularly critical when extrapolating beyond the distribution of the training dataset.
@@
After fully training the conditional diffusion model, it learns the mapping between morphological conditions and corresponding visual features. We call the trained conditional diffusion model GalaxySD which can synthesize galaxies from our dataset and extrapolate to rare or even unseen scenarios, for example, imagine an early-type galaxy exhibiting star-forming characteristics. These hypothetical entities, though rare or even never observed, hold scientific value. Incorporating these synthesized images into downstream machine learning model training pipelines can notably improve model performance on ill-defined problems and may lead to potential unique scientific discoveries.
III Data